# Janus: Resilient and Adaptive Data Transmission for Enabling Timely and Efficient Cross-Facility Scientific Workflows

### v1

@@ 3.2.1. Minimizing Data Transmission Time with a Guaranteed Error Bound on the Received Data
No matter which equation is used to calculate $p$, increasing $m$ has two opposing effects on $E[T_{total}]$: 1) As $N=\frac{S}{(n-m)s}$, a larger $m$ increases the number of FTGs, leading to more fragments transmitted during both the initial transfer and potential retransmissions, thus increasing $E[T_{total}]$. 2) A larger $m$ reduces the probability $p$ of unrecoverable data loss within an FTG, decreasing the likelihood of retransmissions and mitigating the increase in $E[T_{total}]$ caused by a higher $N$. As a result, an optimal value of $m$ exists that minimizes the expected total transmission time $E[T_{total}]$. In practice, since $p$ is relatively small, $E[T_{total}]$ usually converges for $i>50$. Given that $m$ is an integer and $m\leq n/2$ (as the number of parity fragments must not exceed the number of data fragments), finding the optimal $m$ that minimizes $E[T_{total}]$ is computationally straightforward. This can be achieved through an exhaustive search or a lightweight numerical optimization approach.
3.2.2. Minimizing Errors in the Received Data with a Guaranteed Data Transmission Time
In certain scenarios, scientists prioritize receiving data rapidly within a deterministic timeframe, even if it contains some level of error, rather than waiting for a prolonged and unpredictable period to obtain highly accurate data. For example, in scientific visualization, quickly rendering preliminary representations of large datasets allows researchers to identify trends, detect anomalies, or formulate hypotheses, even if the data is imperfect.
To ensure predictable transmission times, retransmissions must be minimized to avoid introducing non-deterministic delays. As a result, fault tolerance relies primarily on parity fragments generated by erasure coding. The key challenge is to determine which levels to transmit and the appropriate number of parity fragments created for the FTG at each level, so that the total transmission time stays within the user-specified timeframe while minimizing errors in the received data. The optimization model presented below is developed to address this challenge.
Assume the size of the $j$-th level of the refactored data is $S_{j}$, and $m_{j}$ parity fragments are generated for this level. Consequently, the number of FTGs required for the first level can be expressed as: $N_{j}=\frac{S_{j}}{(n-m_{j})s}$. If the first $l$ out of $L$ levels of the refactored data are sent without retransmission, the total transmission time can be calculated as:
(9) $\begin{split}T_{total}=&t+\frac{n(N_{1}+N_{2}+\ldots+N_{l})-1}{r}\\ =&t+\frac{n\sum\limits_{j=1}^{l}\frac{S_{j}}{(n-m_{j})s}-1}{r}\end{split}$
Let the user-specified constraint on total transmission time be $\tau$. Given that $20\leq m_{j}\leq n/$ 2, the range of $l$ can be roughly estimated by solving the following inequality:
(10) $t+\frac{n\sum\limits_{j=1}^{l}\frac{S_{j}}{(n-0)s}-1}{r}\leq T_{total}\leq t+\frac{n\sum\limits_{j=1}^{l}\frac{S_{j}}{(n-\frac{n}{2})s}-1}{r}\leq\tau$
@@ 3.2.2. Minimizing Errors in the Received Data with a Guaranteed Data Transmission Time
(11) $\begin{split}E[\varepsilon]=&[1-(1-p_{1})^{\frac{S_{1}}{(n-m_{1})s}}]\varepsilon_{0}+\\ &\sum\limits_{i=2}^{l-1}\prod\limits_{j=1}^{i-1}(1-p_{j})^{\frac{S_{j}}{(n-m_{j})s}}[1-(1-p_{i})^{\frac{S_{i}}{(n-m_{i})s}}]\varepsilon_{i-1}+\\ &\prod\limits_{j=1}^{l}(1-p_{j})^{\frac{S_{j}}{(n-m_{j})s}}\varepsilon_{l}\end{split}$
@@
(12) $\begin{split}\operatorname*{arg\,min}_{m_{j},\forall j\in\{1,\ldots,l\}}&\quad[1-(1-p_{1})^{\frac{S_{1}}{(n-m_{1})s}}]\varepsilon_{0}+\\ &\quad\sum\limits_{i=2}^{l-1}\prod\limits_{j=1}^{i-1}(1-p_{j})^{\frac{S_{j}}{(n-m_{j})s}}[1-(1-p_{i})^{\frac{S_{i}}{(n-m_{i})s}}]\varepsilon_{i-1}+\\ &\quad\prod\limits_{j=1}^{l}(1-p_{j})^{\frac{S_{j}}{(n-m_{j})s}}\varepsilon_{l}\\ \text{s.t.}&\quad t+\frac{n\sum\limits_{j=1}^{l}\frac{S_{j}}{(n-m_{j})s}-1}{r}\leq\tau\end{split}$
This model can be solved using numerical solvers such as SCIP (Bolusani et al., 2024) or other optimization tools capable of handling nonlinear integer programming.
4. Adaptive Data Transfer Protocols
In real-world networks, packet loss rates often fluctuate over time, introducing variability that can impact data transfer performance. Consequently, the fault-tolerance configurations derived from optimization models based on constant packet loss rates may become suboptimal, potentially degrading transmission efficiency or violating the user-specified constraints on total transmission time. To address this challenge, we integrate adaptive mechanisms into the data transfer protocols, which can dynamically adjust fault-tolerance configurations in response to changing packet loss rates, ensuring that the two user requirements outlined in Section 3.2 are consistently satisfied.
4.1. Data Transfer with Guaranteed Error Bound
@@
             foreach $k$ ($k=n-m$) data fragments do
@@
                   Store data and parity fragments of this FTG to $buf$;
@@
                   Wait for $lost\_FTGs\_list$ from Receiver;
                   if $lost\_FTGs\_list$ is empty then
@@
                  foreach $lost\_FTG$ in $lost\_FTGs\_list$ do
                         Obtain fragments of $lost\_FTG$ from $buf$ and transmit;
@@
                         Increment $lost$ by number of lost fragments;
@@
                                     Add $FTG$ to $lost\_FTGs\_list$;
@@
                   Send $lost\_FTGs\_list$ to Sender;
                   if $lost\_FTGs\_list$ is empty then
@@
The sender is designed with multithreading capabilities to manage parity generation and data transmission simultaneously. In particular, the parity generation thread begins by analyzing the user-specified error bound to determine the appropriate number of levels for transmission. It then partitions the data into fragments and calculates the required number of parity fragments ($m$) needed for each FTG. This calculation involves solving the optimization problem defined in Equation 8, which depends on parameters such as the actual packet transmission rate ($r$) and the packet loss rate ($\lambda$) in lost packets per second. Notably, the actual packet transmission rate is determined as the minimum of the parity fragment generation rate ($r_{ec}$) and the network link transmission rate ($r_{link}$) for UDP packet of size $s$. The parity generation thread continuously generates $m$ parity fragments for every $k$ data fragments to form an FTG while monitoring for updates to $\lambda$ from the receiver. If the packet loss rate changes, the thread dynamically recalculates $m$, ensuring the system adapts to current network conditions. Simultaneously, the transmission thread handles the actual sending of fragments in FTGs. Upon completing the transmission of all data fragments, it notifies the receiver and waits for feedback on any lost FTGs. The thread then retransmits the lost FTGs iteratively until list of lost FTGs returned by the receiver is empty.
@@
4.2. Data Transfer with Guaranteed Time
@@
       Input: $\tau$, $[S_{1},S_{2},\ldots,S_{L}]$, $[\varepsilon_{1},\varepsilon_{2},\ldots,\varepsilon_{L}]$, $s$, $t$, $n$, $r_{ec}$, $r_{link}$, $\lambda$
@@
                   Obtain $[m_{1},m_{2},\ldots,m_{l^{\prime}}]$ by solving Equation 12;
@@
                         Store data and parity fragments of this FTG to $buf$;
@@
                         Increment $lost$ by number of lost fragments;
@@
Similar to algorithm 1, the sender in algorithm 2 leverages multithreading to manage parity generation and data transmission concurrently. The parity generation thread begins by calculating the effective packet transmission rate ($r$), determined as the minimum of the parity fragment generation rate ($r_{ec}$) and the network link transmission rate ($r_{link}$). Using this transmission rate, it solves Equation 10 to identify all feasible values for number of levels that can be transmitted within the given deadline $\tau$. If no feasible value exists, the protocol raises an exception, indicating that the specified deadline is too stringent for the current system configuration. Once all feasible values are identified, the parity generation thread refines the selection by determining the optimal number of levels $l$ to transmit and calculating the most effective combination of redundancy parameters $[m_{1},m_{2},\ldots,m_{l}]$ for each level. This is achieved by solving the optimization problem defined in Equation 12, which minimizes the overall error bound ($\varepsilon$) while satisfying the timing and transmission constraints. Each level is then divided into data fragments of size $s$. For the $i$-th level, the parity generation thread generates $m_{i}$ parity fragments for every $k_{i}$ ($k_{i}=n-m_{i}$) data fragments to form an FTG. The parity generation thread is designed to dynamically adapt to changes in network conditions. If the receiver notifies the sender of an updated packet loss rate ($\lambda$), the thread recalculates the redundancy parameters $m_{i}$ by solving Equation 12. However, FTGs that have already been transmitted are excluded from consideration, and the time elapsed since the transmission began is deducted from the remaining deadline $\tau$ when solving Equation 12.
@@
5. Evaluation
Simulation serves as the primary evaluation method for the proposed adaptive data transfer protocols due to its ability to provide a controlled and flexible environment. Unlike real networks, where packet loss rates are unpredictable and difficult to control, simulation allows systematic exploration of protocol behavior under diverse and extreme packet loss conditions. To complement the simulation-based evaluation, we also conducted real network testing focused on validating the functionality of the proposed protocols in practical scenarios.
5.1. Testing Data and Network
The scientific data used in our evaluation was generated by the Nyx cosmology simulation (NYX simulation, 2019) and refactored into a hierarchical representation with four levels using the pMGARD (Liang et al., 2021) library. The sizes of the four levels were $S_{1}=668$ MB, $S_{2}=2.67$ GB, $S_{3}=5.42$ GB, and $S_{4}=17.99$ GB. The relative L-infinity errors for reconstructing the data using the first $i$ levels of the hierarchy, denoted as $\varepsilon_{i}$, were set as follows: $\varepsilon_{1}=0.004$, $\varepsilon_{2}=0.0005$, $\varepsilon_{3}=0.00006$, and $\varepsilon_{4}=0.0000001$. Note that we do not evaluate data refactoring in this work, as it is typically performed during the data generation process and stored on the system where it is produced before any transfer occurs. Evaluations of data refactoring techniques can be found in existing works, such as (Liang et al., 2021; Wan et al., 2023). The testing network was configured with the sender deployed on a workstation at our university and the receiver operating on a virtual machine hosted by CloudLab (CloudLab, 2025), a cloud infrastructure that is part of the National Science Foundation’s NSFCloud program.
5.2. Simulation-Based Evaluation
5.2.1. Design of the Simulation Model
We developed a discrete-event simulation framework using SimPy (SimPy, 2025), a Python library for modeling discrete-event systems. The core simulation models key operations of data transfer protocols. A process models the sender, which continuously generates and transmits packets. For every $k$ data fragments, the sender generates $m$ parity fragments, transmitting each packet every $1/r$ seconds, where $r$ is the transmission rate. Each packet experiences a transmission latency of $t$ seconds. If the packet is not marked as lost by the packet loss process, it is added to a FIFO queue. Another process models the receiver, which retrieves packets from the FIFO queue. The packet loss process simulates losses by generating random time intervals between losses. When a loss event occurs, the packet is marked as lost if the loss event queue is not empty. Afterward, the loss event queue is cleared. A separate process handles control message exchanges between the sender and receiver.
@@ 5.2.1. Design of the Simulation Model
5.2.2. Simulation Parameter Settings
@@
To determine the simulation parameters, we conducted extensive measurements of data transmission over the testing network. First, we measured the transmission latency $t$ by sending 4,096-byte UDP data fragments for 15 minutes and recording the latency of each fragment. These tests were repeated across multiple days to account for temporal variations. The results showed stable UDP transmission latency ranging from 0.0095 to 0.015 seconds, so we set $t=0.01$ seconds for our simulation.
@@
Next, we measured the maximum network transmission rate $r_{link}$ and the parity fragment generation rate $r_{ec}$. Using 4,096-byte packets, we found that $r_{link}=19,144$ packets per second. To determine $r_{ec}$, we used the “liberasurecode” library (Unified Erasure Coding interface for common storage workloads, 2022) to generate parity fragments for FTGs, each consisting of $n=32$ fragments. As $m$ (the number of parity fragments) increased from 1 to 16, $r_{ec}$ decreased from 319,531 to 41,561 fragments per second. Since $r_{ec}$ remained higher than $r_{link}$ even at $m=16$, the actual transmission rate was limited by the network link, $r=r_{link}$. However, in high-bandwidth networks where $r_{link}>r_{ec}$, transmission could be limited by $r_{ec}$.
We determined packet loss rates based on common wide-area network classifications: low, medium, and high. A low loss rate (under 0.1%) corresponds to $\lambda_{low}=r\times 0.1\%=19$ lost packets per second. A medium loss rate (2%) results in $\lambda_{medium}=r\times 2\%=383$ lost packets per second, while a high loss rate (5%) gives $\lambda_{high}=r\times 5\%=957$ lost packets per second. The packet loss generation process uses an exponential distribution with these rate parameters to model the time intervals between consecutive packet loss events.
@@
To model time-varying packet loss rates, we use a hidden Markov model (HMM), as suggested in studies such as (Rodbro et al., 2006; Rossi et al., 2006; Silveira and de Souza e Silva, 2007). The HMM consists of three states—low, medium, and high—each associated with a Gaussian distribution. Specifically, the low state has $\mu=19$, $\sigma=2$, the medium state has $\mu=383$, $\sigma=40$, and the high state has $\mu=957$, $\sigma=100$. Transitions between states follow a continuous-time Markov chain with an exponential holding time distribution, where the rate parameter is 0.04, meaning state transitions occur approximately every 25 seconds. Within each state, the packet loss rate is randomly generated based on the corresponding Gaussian distribution, simulating more realistic, time-varying network conditions.
@@
Finally, protocol-specific parameters are configured using commonly accepted values. For TCP, the retransmission timeout is set to twice the transmission latency, and the duplicate ACK threshold for triggering fast retransmission is set to 3. For the adaptive data transfer protocols, the time window for measuring and updating the packet loss rate is set to 3 seconds, balancing responsiveness to changing network conditions with measurement stability.
5.2.3. Data Transfer under Static Packet Loss Rates

### v2

@@ III-B1 Minimizing Data Transmission Time with a Guaranteed Error Bound on the Received Data
Regardless of how $p$ is calculated, increasing $m$ has two opposing effects on $E[T_{total}]$: first, since $N=\frac{S}{(n-m)s}$, a larger $m$ increases the number of FTGs, resulting in more fragments transmitted during both initial transfer and retransmissions, which raises $E[T_{total}]$; second, a larger $m$ reduces $p$, lowering the likelihood of retransmissions and offsetting the increase caused by a higher $N$. Consequently, an optimal $m$ exists that minimizes $E[T_{total}]$. In practice, with relatively small $p$, $E[T_{total}]$ converges for $i>50$. Since $m$ is an integer and must satisfy $m\leq n/2$ (the number of parity fragments cannot exceed the number of data fragments), determining the optimal $m$ is straightforward using either an exhaustive search or a lightweight numerical optimization.
III-B2 Minimizing Errors in the Received Data with a Guaranteed Data Transmission Time
In some scenarios, scientists prioritize receiving data quickly within a deterministic timeframe, even if it contains some error, rather than waiting for highly accurate data over a prolonged, unpredictable period. For example, in scientific visualization, rapidly rendering preliminary representations of large datasets allows researchers to identify trends, detect anomalies, or formulate hypotheses, even with imperfect data. To ensure predictable transmission times, retransmissions are minimized, relying primarily on parity fragments from erasure coding. The key challenge is determining which levels to transmit and how many parity fragments to generate for each FTG so total transmission time stays within the user-specified limit while minimizing errors. The following optimization model addresses this challenge.
Let the size of the $j$-th level of refactored data be $S_{j}$, with $m_{j}$ parity fragments generated. The number of FTGs for level $j$ is $N_{j}=\frac{S_{j}}{(n-m_{j})s}$. If the first $l$ of $L$ levels are sent without retransmission, the total transmission time is
$\begin{split}T_{total}=&t+\frac{n(N_{1}+N_{2}+\ldots+N_{l})-1}{r}\\ =&t+\frac{n\sum\limits_{j=1}^{l}\frac{S_{j}}{(n-m_{j})s}-1}{r}\end{split}$ (8)
Given a user-specified time constraint $\tau$ and $0\leq m_{j}\leq n/2$, the feasible range of $l$ can be estimated by
$t+\frac{n\sum\limits_{j=1}^{l}\frac{S_{j}}{(n-0)s}-1}{r}\leq T_{total}\leq t+\frac{n\sum\limits_{j=1}^{l}\frac{S_{j}}{(n-\frac{n}{2})s}-1}{r}\leq\tau$ (9)
@@ III-B2 Minimizing Errors in the Received Data with a Guaranteed Data Transmission Time
$\begin{split}E[\varepsilon]=&[1-(1-p_{1})^{\frac{S_{1}}{(n-m_{1})s}}]\varepsilon_{0}+\\ &\sum\limits_{i=2}^{l-1}\prod\limits_{j=1}^{i-1}(1-p_{j})^{\frac{S_{j}}{(n-m_{j})s}}[1-(1-p_{i})^{\frac{S_{i}}{(n-m_{i})s}}]\varepsilon_{i-1}+\\ &\prod\limits_{j=1}^{l}(1-p_{j})^{\frac{S_{j}}{(n-m_{j})s}}\varepsilon_{l}\end{split}$ (10)
@@
$\begin{split}\operatorname*{arg\,min}_{m_{j},\forall j\in\{1,\ldots,l\}}&[1-(1-p_{1})^{\frac{S_{1}}{(n-m_{1})s}}]\varepsilon_{0}+\\ &\sum\limits_{i=2}^{l-1}\prod\limits_{j=1}^{i-1}(1-p_{j})^{\frac{S_{j}}{(n-m_{j})s}}[1-(1-p_{i})^{\frac{S_{i}}{(n-m_{i})s}}]\varepsilon_{i-1}\\ &+\prod\limits_{j=1}^{l}(1-p_{j})^{\frac{S_{j}}{(n-m_{j})s}}\varepsilon_{l}\\ \text{s.t.}&\quad t+\frac{n\sum\limits_{j=1}^{l}\frac{S_{j}}{(n-m_{j})s}-1}{r}\leq\tau\end{split}$ (11)
This model can be solved using numerical solvers such as SCIP [BolusaniEtal2024OO] or other optimization tools capable of handling nonlinear integer programming.
IV Adaptive Data Transfer Protocols
In real-world networks, packet loss rates often fluctuate over time, making fault-tolerance configurations based on constant loss rates suboptimal and potentially degrading transmission efficiency or violating the user-specified constraints on total transmission time. To address this, we incorporate adaptive mechanisms into the data transfer protocols, dynamically adjusting fault-tolerance settings in response to changing loss rates to consistently satisfy the two user requirements outlined in Section III-B.
IV-A Data Transfer with Guaranteed Error Bound
@@
    foreach $k$ (k=nmk=n-m) data fragments do
@@
         Store data and parity fragments of this FTG to bufbuf;
@@
         Wait for lost_FTGs_listlost\_FTGs\_list from Receiver;
       if lost_FTGs_listlost\_FTGs\_list is empty then
@@
       foreach lost_FTGlost\_FTG in lost_FTGs_listlost\_FTGs\_list do
            Obtain fragments of lost_FTGlost\_FTG from bufbuf and transmit;
@@
            Increment lostlost by number of lost fragments;
@@
                  Add FTGFTG to lost_FTGs_listlost\_FTGs\_list;
@@
         Send lost_FTGs_listlost\_FTGs\_list to Sender;
       if lost_FTGs_listlost\_FTGs\_list is empty then
@@
The sender is multithreaded to handle parity generation and data transmission concurrently. The parity generation thread first analyzes the user-specified error bound to determine the levels to transmit, partitions the data into fragments, and calculates the required number of parity fragments ($m$) for each FTG by solving the optimization problem in Equation 7, which depends on the actual packet transmission rate ($r$) and packet loss rate ($\lambda$). The transmission rate $r$ is the minimum of the parity generation rate ($r_{ec}$) and the network link rate ($r_{link}$) for UDP packets of size $s$. The thread continuously generates $m$ parity fragments for every $k$ data fragments while monitoring updates to $\lambda$ from the receiver, dynamically recalculating $m$ as network conditions change. Simultaneously, the transmission thread sends fragments in FTGs, notifies the receiver upon completion, and retransmits lost FTGs iteratively until none remain.
@@
IV-B Data Transfer with Guaranteed Time
@@
 Input: $\tau$, $S_{1},S_{2},\ldots,S_{L}]$, $\varepsilon_{1},\varepsilon_{2},\ldots,\varepsilon_{L}]$, $s$, $t$, $n$, $r_{ec}$, $r_{link}$, $\lambda$
@@
         Obtain $m_{1},m_{2},\ldots,m_{l^{\prime}}]$ by solving Equation 11;
@@
            Store data and parity fragments of this FTG to bufbuf;
@@
            Increment lostlost by number of lost fragments;
@@
The sender in algorithm 2 also uses multithreading to handle parity generation and data transmission concurrently. The parity generation thread first calculates the effective packet transmission rate $r$, the minimum of the parity generation rate ($r_{ec}$) and network link rate ($r_{link}$). Using $r$, it solves Equation 9 to identify feasible numbers of levels that can be transmitted within the deadline $\tau$. If no feasible solution exists, the protocol raises an exception indicating the deadline is too tight. Among feasible options, the thread selects the optimal number of levels $l$ and computes the best redundancy parameters $m_{1},\dots,m_{l}]$ by solving Equation 11, minimizing overall error $\varepsilon$ while meeting timing constraints. Each level is divided into fragments of size $s$, and for level $i$, $m_{i}$ parity fragments are generated for every $k_{i}=n-m_{i}$ data fragments to form an FTG. The thread dynamically adapts to network changes: when the receiver reports an updated packet loss rate $\lambda$, it recalculates $m_{i}$ using Equation 11, excluding already transmitted FTGs and adjusting $\tau$ for elapsed time.
@@
V Evaluation
Simulation is the primary evaluation method for the proposed adaptive data transfer protocols, offering a controlled environment to systematically explore protocol behavior under diverse and extreme packet loss conditions, which are difficult to manage in real networks. To complement this, we also performed real network tests to validate the protocols’ functionality in practical scenarios.
V-A Testing Data and Network
The scientific data for our evaluation was generated by the Nyx cosmology simulation [Nyx] and refactored into a four-level hierarchical representation using the pMGARD [liang2021error] library. The level sizes were $S_{1}=668$ MB, $S_{2}=2.67$ GB, $S_{3}=5.42$ GB, and $S_{4}=17.99$ GB, with relative L-infinity reconstruction errors $\varepsilon_{1}=0.004$, $\varepsilon_{2}=0.0005$, $\varepsilon_{3}=0.00006$, and $\varepsilon_{4}=0.0000001$. We do not evaluate data refactoring here, as it is typically performed during data generation and stored locally; evaluations of data refactoring techniques can be found in existing works [liang2021error, wan2023rapids]. The testing network used two virtual machines on CloudLab [cloudlab], part of the National Science Foundation’s NSFCloud program. The sender VM was on the d430s cluster at Emulab, University of Utah, while the receiver VM was on the c6420 cluster at Clemson University. Data transfers between the sender and receiver traverse the wide-area network.
V-B Simulation-Based Evaluation
V-B1 Design of the Simulation Model
We developed a discrete-event simulation framework using SimPy [simpy], a Python library for modeling discrete-event systems, to capture key operations of data transfer protocols. The sender continuously generates and transmits packets, producing $m$ parity fragments for every $k$ data fragments and sending each packet every $1/r$ seconds, where $r$ is the transmission rate. Each packet experiences a latency of $t$ seconds, and if not lost, it is added to a FIFO queue. The receiver process retrieves packets from this queue. Packet loss is simulated by generating random intervals between loss events; packets are marked lost if the loss queue is nonempty, after which the queue is cleared. A separate process handles control message exchanges between sender and receiver.
@@ V-B1 Design of the Simulation Model
V-B2 Simulation Parameter Settings
@@
To determine the simulation parameters, we conducted a series of detailed measurements over the testing network to characterize its latency, throughput, and packet loss behavior. We first measured the transmission latency $t$ by continuously sending 4,096-byte UDP data fragments for 15 minutes and recording the latency of each fragment. This experiment was repeated across multiple days to capture potential temporal variations in network performance. The observed latencies were highly stable, ranging from 0.029 to 0.039 seconds, and we therefore used the average value $t=0.034$ seconds in our simulations.
@@
Next, we evaluated the maximum network transmission rate, $r_{link}$, and the parity fragment generation rate, $r_{ec}$. Using 4,096-byte packets, we measured a link throughput of $r_{link}=29{,}000$ packets per second. To quantify $r_{ec}$, we used the “liberasurecode” library [liberasurecode] to generate parity fragments for FTGs consisting of $n=32$ fragments. As the number of parity fragments $m$ increased from 1 to 16, the measured $r_{ec}$ decreased from 319,531 to 41,561 fragments per second due to the growing encoding overhead. Because $r_{ec}$ remained higher than $r_{link}$ even at $m=16$, the actual transmission rate was bounded by the network link capacity, i.e., $r=r_{link}$. However, in high-bandwidth environments where $r_{link}>r_{ec}$, the effective throughput would instead be limited by the encoding rate $r_{ec}$.
We then characterized packet loss behavior empirically using the same testing network. From these measurements, we identified three representative packet loss intensity levels: low ($\lambda_{low}=20$ packets per second), medium ($\lambda_{medium}=400$ packets per second), and high ($\lambda_{high}=1000$ packets per second). The packet loss process was modeled as a Poisson process, in which the inter-loss intervals follow an exponential distribution parameterized by the rate $\lambda$, capturing the random and memoryless nature of loss events in large-scale networks.
@@
To simulate time-varying packet loss, we further adopted a hidden Markov model (HMM) [rodbro2006hidden, rossi2006joint, silveira2007predicting] with three latent states—low, medium, and high—corresponding to different congestion conditions. Each state was associated with a Gaussian distribution of packet loss rates: low ($\mu=20$, $\sigma=2$), medium ($\mu=400$, $\sigma=40$), and high ($\mu=1000$, $\sigma=100$). State transitions followed a continuous-time Markov chain with exponentially distributed holding times (rate = 0.04), resulting in an average transition every 25 seconds. Within each state, packet loss events were sampled from the corresponding Gaussian distribution, producing a realistic, temporally correlated pattern of network variability.
@@
Finally, protocol-specific parameters were configured using widely accepted defaults. For TCP, the retransmission timeout (RTO) was set to twice the measured transmission latency, and the duplicate acknowledgment threshold for triggering fast retransmission was set to 3. For the adaptive data transfer protocols, the observation window for estimating and updating the packet loss rate was configured to 3 seconds, providing a balance between responsiveness to rapid network fluctuations and the stability of loss rate estimation.
V-B3 Data Transfer with Guaranteed Error Bound